Fermi National Laboratory


Accelerator Update for April 20, 2005 - April 22, 2005

Wednesday April 20
The day shift began with Operations monitoring store 4104 and the Recycler stash, with the Antiproton Source (Pbar) not stacking due to septa (A:ISEP2V) repairs, and with the MiniBooNE and Meson MCenter experiments taking beam.

Operations terminated store 4104 at 8 AM.

At 8:31 AM, the Main Injector (MI) beam permit tripped off for the first of three times due to a MECAR trip. It reset okay each time. MECAR was learning in after a MI study period.

At 10:24 AM, a Booster RF station (BRF16) tripped off. It reset okay.

Operations established store 4105 at 10:25 AM with an initial luminosity of 110.27E30. Pbar resumed stacking.

A Pbar expert reported at 10:44 AM that he had completed his repair of A:ISEP2V.

At 11:03 AM, Pbar stopped stacking due to another problem with A:ISEP2V. Experts investigated.

Pbar resumed stacking at 1:04 PM. However, A:ISEP2V was still being investigated.

At 2:28 PM, a Booster power supply (EAPS) tripped off. Operators found no bad status. The supply reset okay.

At 2:55 PM, BRF16 tripped off. It reset okay.

At 2:58 PM, EAPS tripped off on an overload. Operators managed to reset the supply five minutes later.

At 3:06 PM, EAPS tripped off again, but with no trip status. Operators informed an expert and then reset the supply.

At 4:25 PM, A:ISEP2V tripped off due to a current fault and would not reset. Operations contacted experts.

An expert reported that the problem with A:ISEP2V might be in the tunnel. Operators put the Recycler into shot setup to stash antiprotons at 5:25 PM and empty the stack.

Operations reported at 5:44 PM that operators interlocked the NuMI and MINOS areas for aperture scans.

At 5:53 PM, BRF16 tripped off. It reset okay.

Recycler completed stashing antiprotons at 6:19 PM.

At 6:39 PM, Operations tried to put Pbar back into the stacking mode, but it didn't work. Experts continue to investigate.

NuMI experts began their aperture scan study around 7:04 PM.

Operators prepared the Pbar tunnel for an access at 7:05 PM.

At 7:55 PM, Operations reported that there was a problem with the NuMI extraction kicker. Operators discovered that kicker experts had worked on this device. Unfortunately, they left equipment hooked up to the kicker that would not allow it to reach its proper voltage. An expert was called in to adjust the kicker.

Pbar experts reported that A:ISEP2V was fixed at 8:37 PM, but that an LCW leak in the tunnel had to be repaired before stacking could resume. The leak was repaired and the techs were out of the tunnel by 9:33 PM.

At 9:41 PM, Operations reported that a Pbar device (D:IB) wouldn't turn on due to a 480 VAC contactor problem. More experts were called in.

Operations established beam to NuMI at 9:46 PM.

At 11:50 PM, BRF16 tripped off. It reset okay.

Thursday April 21
The midnight shift began with Operations monitoring store 4105 and the stash, with Pbar off due to a D:IB 480 VAC contactor problem, with NuMI conducting studies, and with MiniBooNE and MCenter taking beam.

A Pbar expert reported at 12:33 AM that D:IB was back on line.

Pbar resumed stacking at 1:18 AM.

At 7:42 AM, Operations reported that several Booster RF stations tripped off. Operators then noticed that Linac beam had stopped. Operators investigated and discovered no beam coming from the I- Source. Operators reset the smart rack monitor for the source, which resolved the problem.

MI experts began a study at 10:30 AM.

At 10:53 AM, Operations reported that they had lost the ability to control a Debuncher quadrupole power supply. Experts investigated and had the supply back up by 11:39 AM. Pbar resumed stacking.

At 11:27 AM, a MI RF station (MIRF1) tripped off. It reset okay, but then it began tripping off at 11:47 AM and wouldn't stay reset. An expert investigated.

Operations turned off Linac, Booster, and MI at 12:35 PM and prepared the tunnels for accesses.

At 12:58 PM, the Tevatron (TeV) suffered a quench due to a TeV study that had just started.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls for access at 1:18 PM.

Operations established beam to Linac at 2:58 PM. Beam to Booster followed at 3:03 PM.

MI came out of access and had its beam permit by 5:29 PM.

Pbar resumed stacking at 6:16 PM.

CDF came out of access at 6:44 PM.

Operations established beam to MiniBooNE at 8:07 PM.

Operations put the TeV through a dry squeeze at 8:10 PM.

TeV experts began a study period at 8:57 PM.

Friday April 22
The midnight shift began with both stacking and Switchyard (SY) beam off in preparation for shot setup, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 12:10 AM.

Operations established store 4108 at 2:50 AM with an initial luminosity of 69.08E30. Pbar resumed stacking.

MCenter resumed taking beam at 3:23 AM.

At 3:42 AM, Operations reported having trouble with Pbar stacking. Operators investigated and believed that there was a problem with bunch rotation in the Debuncher and MI. Operators called in an expert, but after he arrived at 5:21 AM the problem cleared itself.

At 5:35 AM, the Debuncher bunch rotation problem reappeared.

The Plan for Friday and the Weekend
The plan for today is to stash the Recycler later this morning. There is a potential problem with a TeV sector C2 wet engine. Cryo system techs will monitor the engine and switch the system to JT valve control if the engine fails. Operations will terminate the store during the evening shift. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/22/2005   email Fermilab